OpenStudy (anonymous):

What is the slope of 8x-2y=1?

OpenStudy (anonymous):

2y=8x-1 ===> y =4x-(1/4) ==> the slope = 4

OpenStudy (anonymous):

it should be y=4x-1/2 and 4 is the slope

OpenStudy (anonymous):

raheen-when you said 2y=8x -1, why did the 2 lose the - ? It went to positive?

OpenStudy (anonymous):

good question , it is actully in a shorcut form becase it's as if transfered to the left side and keeping the x term in the right side

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

I mean transferred y to the right and keeping the x in the left :)

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

oh ok so I add 2y to both side instead of subtracting 8x from the left?

OpenStudy (anonymous):



OpenStudy (anonymous):

it could be done in many ways , the important point here is to take y's to a side and put the x's and constatn to the other side , then doing any division to put it in the form: y = m x + b

OpenStudy (anonymous):

8x-2y=1 2y=8x-1 y=4x-(1/2) the slope is 4